# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 7.4. Graph algebras
Let $Q=(Q^{0},Q^{1},r,s)$ be a directed graph. So $Q^{0}$, $Q^{1}$ are countable sets whose elements are interpreted as vertices and edges respectively, and $r,s:Q^{1}\to Q^{0}$ are the range and source maps. For the standard inverse semigroup $S_{Q}$ associated to $Q$ the associated tight groupoid is the standard Deaconu–Renault groupoid $\mathcal{G}_{Q}$ for the one sided shift on the boundary path space of $Q$, see, for instance, [BKM23, Subsection 7.3]. The groupoid $\mathcal{G}_{Q}$ is amenable, Hausdorff and second countable. For any $p\in[1,\infty]$, by [BKM23, Theorem 7.21] and [GL17, Theorem 6.19] the algebra $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S_{Q})$ is naturally isometrically isomorphic to the graph $L^{p}$-operator algebra $F^{p}(Q)$ considered in [BKM23], [CR19], [CMR23], [Phi12]. Therefore for every non-empty $P\subseteq[1,\infty]$ we call
@@
A path in $Q$ is a sequence $\mu=\mu_{1}\mu_{2}\cdots$ (finite or infinite) where $s(\mu_{i})=r(\mu_{i+1})$ for all $i$. We also treat vertices in $Q^{0}$ as paths of length zero. A finite path $\mu=\mu_{1}\mu_{2}\cdots\mu_{n}$ in $Q$ such that $s(\mu_{n})=r(\mu_{1})$ is called a cycle. A vertex $x\in Q^{0}$ is called singular if it is a source or an infinite receiver, that is if $|r^{-1}(v)|\in\{0,\infty\}$. A path in $Q$ is called boundary if either it is infinite or if it starts in a singular vertex. The set of boundary paths $\partial Q$ is naturally uquiped with the topology that makes it hoemomorphic to the tight spectrum of $S_{Q}$ and it is the unit space of $\mathcal{G}_{Q}$. In particular, for any vertex $v\in Q^{0}$, the cylinder $Z_{v}:=\{\mu\in\partial Q:r(\mu)=b\}$ is compact open in $\partial Q$ and so its indicator $1_{Z_{v}}$ is a projection in $C_{0}(\partial Q)\subseteq F^{P}(Q)$. For two vertices $v,w\in Q^{0}$ we write $v\leq w$ if there is a path $\mu$ in $Q$ that ends in $v$ and starts in $w$.
@@
    If every cycle in $Q$ has an entry then every representation of $F^{P}(\mathcal{G})$ (into any Banach algebra) is injective whenever it is non-zero on projections $1_{Z_{v}}\in C_{0}(\partial Q)$, $v\in Q^{0}$.
@@
We apply Corollary C (in the introduction) to $F^{P}(Q)\cong F^{P}(\mathcal{G}_{Q})$. It is well known — see, for instance, [NO19, Propositions 8.2, 8.3] — that $\mathcal{G}_{Q}$ is topologically free if and only if every cycle in $Q$ has an entry, and $\mathcal{G}_{Q}$ is minimal if and only if $Q$ is cofinal. Thus (2) in the assertion follows from Corollary C(2). Assume now that every cycle in $Q$ has an entry. By Corollary C(1) any representation $\pi:F^{P}(Q)\to B$ is injective whenever its restriction to $C_{0}(\partial Q)$ is injective. But $\pi$ is injective on $C_{0}(\partial Q)$ if it is non-zero on projections $1_{Z_{v}}\in C_{0}(\partial Q)$, $v\in Q^{0}$. This can proved using that then $\pi$ is non-zero on projections spaning finite-dimensional algebras turning $C_{0}(\partial Q)$ into an AF-algebra, cf. the proof of [KPR98, Theorem 3.7], or by applying the criterion in the last part of [BKM23, Theorem 6.15] to $S_{Q}$. This shows (1). Assume now $Q$ is cofinal and that every cycle in $Q$ has an entry. It is well known and easy to see that then $\mathcal{G}_{Q}$ is locally contracting if and only if $Q$ contains a cycle. If this holds, then $F^{P}(Q)$ is purely infinite by Theorem D. If $Q$ has no cycles, then $F^{P}(Q)$ is an inductive limit of finite-dimensional graph algebras, which are $\mathbb{F}$-matricial algebras, see [AAM17, Proposition 2.6.20]. This explains (3). ∎
@@ 7.5. Algebras associated to self-similar group actions on graphs
Both the $G$-action and the restriction maps extend uniquely to paths in $Q$, in such a way that $\sigma_{g}(e\mu)=\sigma_{g}(e)\sigma_{g|_{e}}(\mu)$ and $g|_{e\mu}=(g|_{e})|_{\mu}$ for all $g\in G$, $e\in Q^{1}$ and $\mu\in Q^{*}$ with $s(e)=r(\mu)$, see [EP17, Proposition 2.4]. There is a natural inverse semigroup associated to this data, [EP17, Sections 4, 8]. Namely, the set
@@
Let $\mathbb{F}=\mathbb{C}$ and $P=\{2\}$. Then $F^{P}(G,Q)$ is isomorphic to the universal $C^{*}$-algebra $\mathcal{O}_{G,Q}$ introduced in [EP17, Definition 3.2], because $\mathcal{O}_{G,Q}\cong C^{*}(\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q}))=F^{2}_{\operatorname{tight}}(S_{G,Q})$ by [EP17, Corollary 6.4]. This construction can be specialised to give the $C^{*}$-algebras constructed by Katsura [Kat08]. When the graph has only a single vertex the algebra $F^{2}(G,Q)$ becomes the self-similar group $C^{*}$-algebra of Nekrashevych, who also considered its reduced and essential versions $F^{2}_{\operatorname{r}}(G,Q)$ and $F^{2}_{\operatorname{e}}(G,Q)$, see [Nek09, 3.6, 5.4] and [Nek04].
@@
If $G$ is amenable then $F^{P}(G,Q)=F^{P}_{\operatorname{r}}(G,Q)$.
@@
By [EP17, Corollary 10.18] the underlying groupoid is amenable and we may apply [GL17, Theorem 6.19]. ∎
@@
The assumption that $(G,Q)$ is cofinal is equivalent to $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$ being minimal by [EP17, Theorem 13.6]. The remaining assumption on $(G,Q)$ is equivalent to $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$ being effective, see [EP17, Theorem 14.10] and [EP16, Theorem 4.7]. Also under our assumptions $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$ is locally contracting, see [EP17, Theorem 15.1]. Thus since $F^{P}_{\operatorname{e}}(G,Q)$ is an essential Banach algebra for $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$, cf. Corollary 7.26, $F^{P}_{\operatorname{e}}(G,Q)$ is simple purely infinite by Theorem 6.13. Every $g\in G$ admits at most finitely many minimal strongly fixed paths if and only if $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$ is Hausdorff, see [EP17, Theorem 12.2]. Thus then $F^{P}_{\operatorname{e}}(G,Q)=F^{P}_{\operatorname{r}}(G,Q)$. If [CEPSS19, (5.9)] holds, then $S_{G,Q}$ satisfies condition (S) mentioned in Remark 7.34, and so then $F^{P}_{\operatorname{r}}(G,Q)$ is an essential Banach algebra. ∎
@@
Let $A$ and $B$ be two $N\times N$ matrices with integer entries. Katsura [Kat08] associated to them a $C^{*}$-algebra, and Exel–Pardo noticed in [EP17, Example 3.4] that there is a natural self-similar action $(\mathbb{Z},Q_{A})$ associated to $A$ and $B$ that yield the same algebra. Thus for any non-empty $P\subseteq[1,\infty]$, one could view $F^{P}(\mathbb{Z},Q_{A})=F^{P}_{\operatorname{r}}(\mathbb{Z},Q_{A})$ as $L^{P}$-versions of Katsura algebras. In particular, if the underlying groupoid is Hausdorff, which is characterized in terms of $A$ and $B$ in [EP17, Theorem 18.6], then $F^{P}(\mathbb{Z},Q_{A})$ is simple (and purely infinite) if and only if the matrix $A$ is irreducible, see [EP17, Theorem 18.12]. An example of $A$ and $B$ such that the underlying groupoid is non-Hausdorff but satisfies [CEPSS19, (5.9)] is analyzed in [CEPSS19, Subsection 5.4]. Thus Theorem 7.42 applied to this example gives that $F^{P}(\mathbb{Z},Q_{A})$ is simple and purely infinite.
@@
Let $X$ be a set with at least two elements. Self-similar groups $(G,X)$, cf. [Nek04], [Nek09], are equivalent to self-similar group actions $(G,Q_{X})$ on the graph $Q_{X}$ that has only one vertex, the set of edges $E^{1}=X$, and such that the extended of $G$ on finite paths $X^{*}$ is faithful. Then the associated groupoid is necessarily minimal and topologically free, cf.  [EP17, Corollary 14.14], and so the first part of Theorem 7.42 applies. Thus for any self-similar group $(G,X)$ and any non-empty $P\subseteq[1,\infty]$ the associated essential algebra $F^{P}_{\operatorname{e}}(G,X):=F^{P}_{\operatorname{e}}(G,Q_{X})$ is always purely infinite and simple. If $(G,X)$ is $\omega$-faithful in the sense of [CEPSS19, Definition 5.20], then also [CEPSS19, (5.9)] holds, and so $F^{P}_{\operatorname{r}}(G,X):=F^{P}_{\operatorname{r}}(G,Q_{X})$ is also simple purely infinite in this case. The second part of Theorem 7.42 does not apply to the self-similar action $(G,X)$ modeling the Grigorchuk group. Nevertheless the proof of [CEPSS19, Theorem 5.22] shows that $F^{P}(G,X):=F^{P}(G,Q_{X})$ is simple purely infinite in this case as well.

### v3

@@ 7.4. Graph algebras
Let $Q=(Q^{0},Q^{1},r,s)$ be a directed graph. So $Q^{0}$, $Q^{1}$ are countable sets whose elements are interpreted as vertices and edges respectively, and $r,s:Q^{1}\to Q^{0}$ are the range and source maps. For the standard inverse semigroup $S_{Q}$ associated to $Q$, the associated tight groupoid is the standard Deaconu–Renault groupoid $\mathcal{G}_{Q}$ for the one sided shift on the boundary path space of $Q$, see, for instance, [BKM25, Subsection 6.1]. The groupoid $\mathcal{G}_{Q}$ is amenable, Hausdorff and second countable. For any $p\in[1,\infty]$, by Proposition 7.23, [BKM25, Corollary 6.25] and [GL17, Theorem 6.19], the algebra $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S_{Q})$ is naturally isometrically isomorphic to the graph $L^{p}$-operator algebra $F^{p}(Q)$ considered in [BKM25, Definition 6.22], [CR19], [CMR23], [Phi12]. Therefore, for every non-empty $P\subseteq[1,\infty]$, we call
@@
A path in $Q$ is a sequence $\mu=\mu_{1}\mu_{2}\cdots$ (finite or infinite) where $s(\mu_{i})=r(\mu_{i+1})$ for all $i$. We also treat vertices in $Q^{0}$ as paths of length zero. A finite path $\mu=\mu_{1}\mu_{2}\cdots\mu_{n}$ in $Q$ such that $s(\mu_{n})=r(\mu_{1})$ is called a cycle. A vertex $x\in Q^{0}$ is called singular if it is a source or an infinite receiver, that is, if $|r^{-1}(v)|\in\{0,\infty\}$. A path in $Q$ is called boundary if either it is infinite or if it starts in a singular vertex. The set of boundary paths $\partial Q$ is naturally equipped with the topology that makes it homemomorphic to the tight spectrum of $S_{Q}$ and it is the unit space of $\mathcal{G}_{Q}$. In particular, for any vertex $v\in Q^{0}$, the cylinder $Z_{v}:=\{\mu\in\partial Q:r(\mu)=b\}$ is compact open in $\partial Q$ and so its indicator $1_{Z_{v}}$ is a projection in $C_{0}(\partial Q)\subseteq F^{P}(Q)$. For two vertices $v,w\in Q^{0}$, we write $v\leq w$ if there is a path $\mu$ in $Q$ that ends in $v$ and starts in $w$.
@@
    If every cycle in $Q$ has an entry, then every representation of $F^{P}(\mathcal{G})$ (into any Banach algebra) is injective whenever it is non-zero on projections $1_{Z_{v}}\in C_{0}(\partial Q)$, $v\in Q^{0}$.
@@
We apply Corollary C from the introduction to $F^{P}(Q)\cong F^{P}(\mathcal{G}_{Q})$. It is well known, see, for instance, [NO19, Propositions 8.2, 8.3], that $\mathcal{G}_{Q}$ is topologically free if and only if every cycle in $Q$ has an entry, and $\mathcal{G}_{Q}$ is minimal if and only if $Q$ is cofinal. Thus (2) in the assertion follows from Corollary C(2). Assume now that every cycle in $Q$ has an entry. By Corollary C(1), any representation $\pi:F^{P}(Q)\to B$ is injective whenever its restriction to $C_{0}(\partial Q)$ is injective. But $\pi$ is injective on $C_{0}(\partial Q)$ if it is non-zero on projections $1_{Z_{v}}\in C_{0}(\partial Q)$, $v\in Q^{0}$. This can be proved using that $\pi$ is non-zero on projections spanning finite-dimensional algebras turning $C_{0}(\partial Q)$ into an AF-algebra, cf. the proof of [KPR98, Theorem 3.7], or by applying the criterion in the last part of [BKM25, Theorem 6.15] to $S_{Q}$. This shows (1). Assume now $Q$ is cofinal and that every cycle in $Q$ has an entry. It is well known and easy to see that then $\mathcal{G}_{Q}$ is locally contracting if and only if $Q$ contains a cycle. If this holds, then $F^{P}(Q)$ is purely infinite by Theorem D. If $Q$ has no cycles, then $F^{P}(Q)$ is an inductive limit of finite-dimensional graph algebras, which are $\mathbb{F}$-matricial algebras, see [AAM17, Proposition 2.6.20]. This proves (3). ∎
@@ 7.5. Algebras associated to self-similar group actions on graphs
Both the $G$-action and the restriction maps extend uniquely to paths in $Q$, in such a way that $\sigma_{g}(e\mu)=\sigma_{g}(e)\sigma_{g|_{e}}(\mu)$ and $g|_{e\mu}=(g|_{e})|_{\mu}$ for all $g\in G$, $e\in Q^{1}$ and $\mu\in Q^{*}$ with $s(e)=r(\mu)$, see [EP17, Proposition 2.4]. There is a natural inverse semigroup associated to these data, [EP17, Sections 4, 8]. Namely, the set
@@
Let $\mathbb{F}=\mathbb{C}$ and $P=\{2\}$. Then $F^{P}(G,Q)$ is isomorphic to the universal $C^{*}$-algebra $\mathcal{O}_{G,Q}$ introduced in [EP17, Definition 3.2], because $\mathcal{O}_{G,Q}\cong C^{*}(\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q}))=F^{2}_{\operatorname{tight}}(S_{G,Q})$ by [EP17, Corollary 6.4]. This construction can be specialized to give the $C^{*}$-algebras constructed by Katsura [Kat08]. When the graph has only a single vertex, the algebra $F^{2}(G,Q)$ becomes the self-similar group $C^{*}$-algebra of Nekrashevych, who also considered its reduced and essential versions $F^{2}_{\operatorname{r}}(G,Q)$ and $F^{2}_{\operatorname{e}}(G,Q)$, see [Nek09, 3.6, 5.4] and [Nek04].
@@
If $G$ is amenable, then $F^{P}(G,Q)=F^{P}_{\operatorname{r}}(G,Q)$.
@@
By [EP17, Corollary 10.18], the underlying groupoid is amenable and we may apply [GL17, Theorem 6.19]. ∎
@@
The assumption that $(G,Q)$ is cofinal is equivalent to $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$ being minimal by [EP17, Theorem 13.6]. The remaining assumption on $(G,Q)$ is equivalent to $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$ being effective, see [EP17, Theorem 14.10] and [EP16, Theorem 4.7]. Also, under our assumptions, $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$ is locally contracting, see [EP17, Theorem 15.1]. Thus since $F^{P}_{\operatorname{e}}(G,Q)$ is an essential Banach algebra for $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$, cf. Corollary 7.26, $F^{P}_{\operatorname{e}}(G,Q)$ is simple purely infinite by Theorem 6.13. Every $g\in G$ admits at most finitely many minimal strongly fixed paths if and only if $\mathcal{G}_{{\operatorname{tight}}}(S_{G,Q})$ is Hausdorff, see [EP17, Theorem 12.2]. Thus then $F^{P}_{\operatorname{e}}(G,Q)=F^{P}_{\operatorname{r}}(G,Q)$. If [CEPSS19, (5.9)] holds, then $S_{G,Q}$ satisfies condition (S) mentioned in Remark 7.34, and so then $F^{P}_{\operatorname{r}}(G,Q)$ is an essential Banach algebra. ∎
@@
Let $A$ and $B$ be two $N\times N$ matrices with integer entries. Katsura [Kat08] associated to them a $C^{*}$-algebra, and Exel–Pardo noticed in [EP17, Example 3.4] that there is a natural self-similar action $(\mathbb{Z},Q_{A})$ associated to $A$ and $B$ that yield the same algebra. Thus, for any non-empty $P\subseteq[1,\infty]$, one could view $F^{P}(\mathbb{Z},Q_{A})=F^{P}_{\operatorname{r}}(\mathbb{Z},Q_{A})$ as $L^{P}$-versions of Katsura algebras. In particular, if the underlying groupoid is Hausdorff, which is characterized in terms of $A$ and $B$ in [EP17, Theorem 18.6], then $F^{P}(\mathbb{Z},Q_{A})$ is simple (and purely infinite) if and only if the matrix $A$ is irreducible, see [EP17, Theorem 18.12]. An example of $A$ and $B$ such that the underlying groupoid is non-Hausdorff, but satisfies [CEPSS19, (5.9)], is analyzed in [CEPSS19, Subsection 5.4]. Thus Theorem 7.42 applied to this example gives that $F^{P}(\mathbb{Z},Q_{A})$ is simple and purely infinite.
@@
Let $X$ be a set with at least two elements. Self-similar groups $(G,X)$, cf. [Nek04], [Nek09], are equivalent to self-similar group actions $(G,Q_{X})$ on the graph $Q_{X}$ that has only one vertex, the set of edges $E^{1}=X$, and such that the extended action of $G$ on the set of finite paths $X^{*}$ is faithful. Then the associated groupoid is necessarily minimal and topologically free, cf. [EP17, Corollary 14.14], and so the first part of Theorem 7.42 applies. Thus, for any self-similar group $(G,X)$ and any non-empty $P\subseteq[1,\infty]$, the associated essential algebra $F^{P}_{\operatorname{e}}(G,X):=F^{P}_{\operatorname{e}}(G,Q_{X})$ is always purely infinite and simple. If $(G,X)$ is $\omega$-faithful in the sense of [CEPSS19, Definition 5.20], then also [CEPSS19, (5.9)] holds, and so $F^{P}_{\operatorname{r}}(G,X):=F^{P}_{\operatorname{r}}(G,Q_{X})$ is also simple purely infinite in this case. The second part of Theorem 7.42 does not apply to the self-similar action $(G,X)$ modeling the Grigorchuk group. Nevertheless, the proof of [CEPSS19, Theorem 5.22] shows that $F^{P}(G,X):=F^{P}(G,Q_{X})$ is simple purely infinite in this case as well.